OpenStudy (anonymous):

What is the approximate perimeter of a semicircle with a radius of 9 mm? Use π ⁢ = 3.14. Express your final answer to the nearest hundredth. I THINK IT IS 56.52

OpenStudy (anonymous):

@bohotness @HazelLuv99 @jagr2713 @Anon101 @whydoihavetosignup1 @Quan99 @outdoorgirl12 @ObeyRealz @Kkurayuma @kmorgan @lizz123 @LittleRedRiding @HELPMEPLZ!! @Grace808 @GotThemAnswers @Gokuporter @freegirl112 @magepker728 @mathkid911 @Unanimoose @undeadknight26 @

OpenStudy (anonymous):

= 2(3 + 6) = 2(9) = 18 Answer (1): 18 mm assuming it's a rectangle ---------------- = (12 * 2 * 22/7)/2 + 2(12) = (528/7)/2 + 24 = 264/7 + 168/7 = 432/7 or 61 5/7 Answer (2): 61 5/7 centimeters OR 61.7 1/7 centimeters

OpenStudy (anonymous):

= 2(3 + 6) = 2(9) = 18 Answer (1): 18 mm assuming it's a rectangle ---------------- = (12 * 2 * 22/7)/2 + 2(12) = (528/7)/2 + 24 = 264/7 + 168/7 = 432/7 or 61 5/7 Answer (2): 61 5/7 centimeters OR 61.7 1/7 centimeters

OpenStudy (anonymous):

its not perimeter its circumference

jagr2713 (jagr2713):

Please avoid Mass tagging

OpenStudy (anonymous):

= 2(3 + 6) = 2(9) = 18 Answer (1): 18 mm assuming it's a rectangle ---------------- = (12 * 2 * 22/7)/2 + 2(12) = (528/7)/2 + 24 = 264/7 + 168/7 = 432/7 or 61 5/7 Answer (2): 61 5/7 centimeters OR 61.7 1/7 centimeters

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

@Unanimoose what do u think

OpenStudy (anonymous):

and in order to find the circumference you want to try 2piR

OpenStudy (unanimoose):

Yeah.

OpenStudy (anonymous):

So 2 times pi and the radius

OpenStudy (anonymous):

it is 61.7

OpenStudy (anonymous):

so the radius would be nine

OpenStudy (anonymous):

is it @Unanimoose and if so, how

OpenStudy (anonymous):

right from here multiply pi and the radius then multiply by two

OpenStudy (anonymous):

61.7

OpenStudy (anonymous):

i mean

OpenStudy (anonymous):

okay

OpenStudy (unanimoose):

\[C = 9^2\pi\]

OpenStudy (anonymous):

ohh, so 9

OpenStudy (anonymous):

???

OpenStudy (anonymous):

hhhhhmmmm...

OpenStudy (anonymous):

i got 28.26 lol

OpenStudy (anonymous):

yeah, is it 28.26

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ya i am right yay i am impressed me 6 grader i did it lol

OpenStudy (unanimoose):

\[C = 81\pi\] \[C = 254.\]

OpenStudy (anonymous):

?????

OpenStudy (unanimoose):

254/2 = 127.17

OpenStudy (anonymous):

weird how u get that

OpenStudy (unanimoose):

Um yeah Idk if i did it wrong or right.

OpenStudy (anonymous):

no

OpenStudy (unanimoose):

Because a semicircle you just find the circumference of a regualr circle then half.

OpenStudy (unanimoose):

cuz a semicircle is jsut half a circle

OpenStudy (anonymous):

u did that wrong, cause I can not type it all in

OpenStudy (anonymous):

Did I do it right

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i just pretended the semi circle was full circle found circumference and divided that by two

OpenStudy (anonymous):

i think you found the area of the circle

OpenStudy (anonymous):

it is 28.26, right??

OpenStudy (unanimoose):

Oh i squared it oops.

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

if you used the formula pi time Radius squared you might get your answer that the area of circle

OpenStudy (anonymous):

tats ok

OpenStudy (anonymous):

lol

OpenStudy (unanimoose):

But tyes it's 28.26.

OpenStudy (unanimoose):

yes*

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

up could you press best response please??

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thx guys

OpenStudy (anonymous):

bye up

OpenStudy (anonymous):

I already did, I will tag you again in a different question though

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hold on a sec...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So did I do it right I hope I did

OpenStudy (anonymous):

I do not think u did :(

OpenStudy (anonymous):

yes all you have to do is use the right formula and then everything is accomplishable

OpenStudy (unanimoose):

unless you use the formula wrong lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Awe man sorry

OpenStudy (anonymous):

46.21?

OpenStudy (anonymous):

28.26